OpenStudy (anonymous):

find an equation of the line that has a slope -2 and passes through the point 10,-9

OpenStudy (anonymous):

y-y1= mx(x-x1)

OpenStudy (amistre64):

do you know the equation of a line that passes thru the origin (0,0) that has a slope of -2?

OpenStudy (anonymous):

write an equation in slope intercept form that has points (5,6) (6,8)

OpenStudy (amistre64):

you seem to just want to use this site as a free answering service ...

OpenStudy (anonymous):

no I`m just stuck on these two

OpenStudy (amistre64):

ok, then can you show us what youve tried so far? its prolly something minor that you may be overlooking

OpenStudy (anonymous):

is the formula y+mx+b=x

OpenStudy (amistre64):

almost: y = mx+b , looks better to me

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

your welcome, anything else?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

k, good luck :)

OpenStudy (anonymous):

write the equation of the line that passes through the points (5,6) and (6,8). Rewrite in slope intercept form

OpenStudy (anonymous):

someone help me please

OpenStudy (anonymous):

You need the equation y = mx + b right? First, we find what m is. I'm sure you know m = (y2 - y1)/(x2 - x1) = (8 - 6)/( 6 - 5) = 2 ( =m) So pick one of the points say (x,y) = (5,6) and plug in y = mx + b to find b That is 6 = 2(5) + b So that 6 = 10 + b and b = -4 Finally, the equation is y = 2x - 4 Hope this helps!